# Universality Diagram of Phase Transitions in Long-range Statistical Systems

### v1

@@ V Long-range O($n$) spin model
Figure 2: Universality diagram for the O($n$) models in the $(d,\sigma)$ plane. Six regimes are identified: the no finite-temperature phase transition region (I), the short-range universality regimes (II and V), and the long-range regimes (III–VI) with distinct scaling behaviors. The boundary at $\sigma=2$ separates the SR and LR universality classes, while $\sigma=1$ marks a crossover between two nonclassical LR regimes.
@@
where $\mathbf{S}_{i}$ is an $n$-component unit vector at site $i$, $N=L^{d}$ is the total number of sites, and $r_{ij}$ denotes the minimal-image distance on the torus [frenkel_understanding_2002, christiansen_phase_2019, agrawal_kinetics_2021]. Here, $n=1,2,3$ corresponds to the Ising, XY, and Heisenberg models, respectively. In contrast to percolation, where the order parameter is purely geometric, O($n$) models possess internal spin degrees of freedom. The competition between spatial dimension, interaction range, and internal symmetry leads to rich critical behavior.
@@
In the SR limit $\sigma\to\infty$, the O($n$) models are governed by a $\phi^{4}$ field theory with upper critical dimension $d_{c}=4$. For $d=1$, all O($n$) models have $T_{c}=0$. In $d=2$, the Mermin–Wagner theorem forbids spontaneous breaking of continuous symmetry at any finite temperature, so the nature of the transition depends on $n$: the Ising model undergoes a continuous transition into a long-range ordered phase, the XY model exhibits a BKT transition into a quasi-long-range ordered phase, and the Heisenberg model has no finite-temperature transition but shows asymptotically free behavior at low temperatures [mermin1966, kosterlitz_critical_1974, yao_asymptotic_2025]. For $d>d_{c}$, the SR Gaussian fixed point determines the critical scaling with exponents $y_{t}^{\rm GF}=2$ and $y_{h}^{\rm GF}=\frac{2+d}{2}$. In finite systems with PBCs, complete-graph asymptotics with $(y_{t}^{\rm CG},y_{h}^{\rm CG})=(\frac{d}{2},\frac{3d}{4})$ are required to describe the finite-size scaling fully, as discussed in Sec. II.
@@
When long-range interactions are introduced, the nature of the transition can change substantially, in ways that are parallel to LR-percolation but modified by the internal symmetry. In $d=1$, where SR Ising, XY, and Heisenberg models have no finite-temperature transition, sufficiently slowly decaying interactions ($\sigma\leq 1$) induce nontrivial criticality, and a BKT-like transition has been reported at $(d,\sigma)=(1,1)$ in the LR-Ising and LR-XY models [dysonExistencePhasetransitionOnedimensional1969, takamoto_critical_2010]. For $2\leq d<4$, numerical and field-theoretical analyses indicate three regimes as $\sigma$ is varied. For $\sigma>2$, the LR tail is irrelevant, and SR universality is retained. In a nonclassical regime $d/2<\sigma\leq 2$, the critical exponents become nontrivial functions of $\sigma$. Finally, for $\sigma\leq d/2$, the LR-GFP controls the transition, with exponents given by dimensional analysis as $(y_{t}^{\rm GF},y_{h}^{\rm GF})=(\sigma,(d+\sigma)/2)$. Extensive numerical evidence in $d=2$ for LR-Ising, XY, and Heisenberg systems supports a sharp SR–LR crossover at $\sigma_{*}=2$ [xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange]. In particular, the point $(d,\sigma)=(2,2)$ already differs from the SR case: 2D LR-XY and LR-Heisenberg models exhibit second-order transitions, and the 2D LR-Ising exponents show detectable deviations from their SR values [xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange, LRHeisenberg].
@@
Unlike the LR-percolation model, rigorous results for LR-O($n$) models in the low-dimensional nonclassical regime are lacking, but heuristic guidance can be obtained from LR-SRW. The LR-SRW effectively describes the Goldstone modes in the ordered phase of LR spin systems [LRHeisenberg]. Specifically, when incorporating a fixed number of steps of $O(L^{d})$ to account for the extensivity of a finite system, the LR-SRW correctly reproduces the finite-size scaling of observables in the LRO phase, such as magnetic susceptibility. Notably, this framework succeeds even at the marginal case $\sigma=2$, where it correctly captures the existence of LRO and the logarithmic scaling of correlations that the standard Gaussian Free Field description fails to predict [xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange, LRHeisenberg]. The LR-SRW also provides a simple geometric explanation for the no-transition regime, in analogy to the argument in the LR-percolation case. Given that the LR-SRW captures these essential physical features, the transition of the random walk from superdiffusive to hyperballistic at $\sigma=1$ leads us to conjecture a similar change in the critical behavior of O($n$) models at this boundary.
@@
Combining SR field theory, LR-GFP analysis, numerical results, and Lévy-flight arguments yields the universality diagram shown in Fig. 2. The overall topology in the $(d,\sigma)$ plane mirrors that of LR-percolation: we again distinguish six regimes, labeled $T_{c}=0$, SR-LD, LR-LD-A, LR-LD-B, SR-HD, and LR-HD. The main structural differences are that the SR upper critical dimension is now $d_{c}=4$ (instead of $6$ for percolation) and that the boundary between LR low-dimensional and LR high-dimensional regimes is given by $\sigma=d/2$. Since several sectors closely parallel the percolation case discussed in Sec. IV, here we only summarize the main regimes and emphasize features specific to O($n$) models.
@@
  • Regime VI (LR-HD): $d>2\sigma$ and $\sigma<2$. The LR Gaussian fixed point with $(y^{\rm GF}_{t},y^{\rm GF}_{h})=\left(\sigma,\frac{d+\sigma}{2}\right)$ controls the critical scaling in the thermodynamic limit, while the CG exponents $(y_{t}^{\rm CG},y_{h}^{\rm CG})=(\frac{d}{2},\frac{3d}{4})$ govern the finite-size scaling of global observables, as in the LR-HD regime of LR-percolation.
@@
  • $\sigma=d$ for $d<2$: This line marks the boundary of the $T_{c}=0$ regime, based on LR SRW arguments. As in LR-percolation, the nature of phase transitions along this line remains unknown, in general. The point $(d,\sigma)=(1,1)$ is a notable example in which LR-Ising and LR-XY models exhibit BKT-like transitions.
@@
  • $\sigma=2$ for $2<d<4$: The line $\sigma=2$ separates SR and LR universality sectors but remains LR-dominated rather than SR-like. Numerical studies at $d=2$ show that at $(d,\sigma)=(2,2)$ the XY and Heisenberg models exhibit second-order transitions and the Ising exponents deviate from their SR values [xiao_two-dimensional_2024, yao2025nonclassicalregimetwodimensionallongrange, LRHeisenberg]. More generally, discontinuities in exponents and universal ratios as functions of $\sigma$ are expected when crossing this line within the LR-LD-A sector.
@@
  • $d=2\sigma$, $\sigma<2$: The LR analogue of the SR upper critical dimension. Here, the LR-GFP becomes marginal, and logarithmic corrections are expected, similarly to the role played by $d=4$ in the SR case.
@@ VI Long-range FK-Ising model
In addition to the standard spin representation, the Ising model admits a geometric formulation via the FK random-cluster representation [grimmett2006], with the partition function defined in Eq. (1). In this framework, percolation and the Ising model share a common geometric language expressed through FK clusters.
Recent studies on the FK representation of the Ising model indicate the existence of two distinct upper critical dimensions [fang_geometric_2022, fang_geometric_2023, wiese_two_2024, engelenburg_one-arm_2025]. The standard thermodynamic upper critical dimension, $d_{c}=4$, marks the onset of mean-field critical behaviors. On the other hand, the cluster geometry exhibits a second, geometric upper critical dimension at $d_{p}=6$, above which the critical clusters, except for the largest one, become percolation-like. Following Ref. [fang_geometric_2022], it is convenient to parameterize the scaling of the largest and second-largest FK clusters by four fractal dimensions $(d_{\mathrm{L}1},d_{\mathrm{F}1},d_{\mathrm{L}2},d_{\mathrm{F}2})$ defined through
@@
where $C_{1}$ and $C_{2}$ are the sizes of the largest and second-largest FK clusters, $R_{1}$ and $R_{2}$ are their unwrapped gyration radii, and $L$ is the linear system size. Here, $d_{\mathrm{L}i}$ denotes the finite-size fractal dimension, characterizing the scaling with $L$, while $d_{\mathrm{F}i}$ represents the thermodynamic fractal dimension, relating the cluster size to its unwrapped gyration radius. For $4<d<6$, the FK clusters are geometrically Ising-like, described by $d_{\mathrm{F}1}=d_{\mathrm{L}1}=\frac{3d}{4}$ and $d_{\mathrm{F}2}=d_{\mathrm{L}2}=\frac{2+d}{2}$. The scaling of the largest cluster size is consistent with asymptotics derived rigorously on the CG Ising model [luczak_phase_2006, bollobas_random-cluster_1996], while the scaling of the second-largest clusters is inferred from numerical results and found to be consistent with the GFP asymptotics [fang_geometric_2022, fang_geometric_2023]. The winding number of the clusters and the number of spanning clusters are both of order $O(1)$. For $d\geq 6$, the finite-size fractal dimensions $d_{\mathrm{L}1}$ and $d_{\mathrm{L}2}$ remain the same as in the lower-dimensional case, but the thermodynamic fractal dimensions of clusters become $d_{\mathrm{F}1}=9/2$ and $d_{\mathrm{F}2}=4$, independent of dimensionality. Notably, $d_{\mathrm{F}2}=4$ corresponds to high-dimensional percolation; thus, the critical clusters, except for the largest one, show percolation-like behavior.
Moreover, the cluster number density $n(s,L)$ differs in the two regimes. Generally, one expects $n(s,L)\sim s^{-\tau}\tilde{n}(s/s_{c})$, where $s_{c}$ is a cutoff size close to the size of the largest cluster, scaling as $s_{c}\sim L^{d_{\mathrm{L}1}}$. The Fisher exponent $\tau$ is believed to satisfy the hyperscaling relation $\tau=1+d/d_{\mathrm{L}1}$, which has been confirmed for percolation in various dimensions and for the random-cluster model for $d<4$ [hou_geometric_2019]. For $4<d<6$, the Fisher exponent follows the GFP prediction as $\tau=1+d/y_{h}^{\rm GF}$ with $y_{h}^{\rm GF}=\frac{2+d}{2}$. However, for $d\geq 6$, the hyperscaling breaks down and one finds $\tau=5/2$, consistent with high-$d$ percolation. These differences in geometric scaling relations further lead to distinctive topological properties in the two regimes [fang_geometric_2022, fang_geometric_2023].
@@
The same mechanism extends naturally to the LR-HD sector for $\sigma<\min(2,d/2)$, where the FSS of thermodynamic properties is governed by the LR Gaussian fixed point with $(y_{t}^{\rm GF},y_{h}^{\rm GF})=(\sigma,(d+\sigma)/2)$ and CG-Ising asymptotics. We conjecture that there exists a $\sigma$-dependent geometric upper critical dimension at $d_{p}=3\sigma$ within the LR mean-field regime, which divides it into two new regimes, LR-HD-A and LR-HD-B, as shown in Fig. 3.
  • Regime VI (LR-HD-A): $2\sigma<d<3\sigma$ with $\sigma<2$. Analogous to the SR-HD-A regime, the largest cluster follows the CG-Ising asymptotics, and the second largest clusters follow the LR-GFP scaling, i.e., $d_{\mathrm{F}1}=d_{\mathrm{L}1}=\frac{3d}{4}$ and $d_{\mathrm{F}2}=d_{\mathrm{L}2}=\frac{\sigma+d}{2}$. Both clusters scale linearly with the system size, as $R_{1}\sim R_{2}\sim L$. By analogy, we expect small clusters to have fractal dimension $d_{\mathrm{L}2}=(d+\sigma)/2$, and the Fisher exponents are $\tau=1+2d/(d+\sigma)$. Therefore, the number of spanning clusters remains finite in the thermodynamic limit.
  • Regime VIII (LR-HD-B): $d\geq 3\sigma$ and $\sigma<2$. In this regime, the critical clusters, except for the largest one, are expected to exhibit LR-percolation-like behavior, i.e., $d_{\mathrm{F}2}=2\sigma$, while the largest cluster is conjectured to have $d_{\mathrm{F}1}=\frac{9}{4}\sigma$. A notable difference from the SR-HD-B regime is that both $d_{\mathrm{F}1}$ and $d_{\mathrm{F}2}$ are $\sigma$-dependent, and the latter is consistent with those of the LR-percolation as listed in Table 2. The finite-size fractal dimensions $d_{\mathrm{L}1},d_{\mathrm{L}2}$ remain the same as in LR-HD-A. The unwrapped gyration radii of clusters are then given by,
@@
    Similar to the SR-HD-B regime, these clusters wind extensively across the periodic boundaries. Additionally, based on the insight of the CG percolation model [huang2018], the Fisher exponent is also expected to be the same as the high-$d$ percolation case, with $\tau=5/2$. The number of spanning clusters then grows as $N_{s}\sim L^{d-3\sigma}$.
@@
This geometric upper critical dimension can be further corroborated by the structure of scaling windows near the critical point. As detailed in Ref. [fang_geometric_2022, fang_geometric_2023] for the SR case, for $d>4$, the FK-Ising model exhibits two critical scaling windows. The first one is of size $O(L^{-d/2})$, consistent with the CG-Ising exponent $y_{t}^{\rm CG}=\frac{d}{2}$. In this interval, the FSS behaviors are identical to those at the critical point. However, a second scaling window exists in the high-T regime, where the FSS differs. For $4<d<6$, the size of the second window is controlled by the GFP exponent $y_{t}=2$. In this scaling window, the various observables follow the GFP scaling. For $d>6$, a percolation scaling window of size $O(L^{-d/3})$ emerges, where the scaling of all clusters is percolation-like. Similar scaling behaviors are expected in the LR-FK-Ising model and are summarized below.
@@
SR-HD-A $\frac{3}{4}d$ $\frac{3}{4}d$ $\frac{d+2}{2}$ $\frac{d+2}{2}$ $\frac{2d}{d+2}$
SR-HD-B $\frac{3}{4}d$ $\frac{9}{2}$ $\frac{d+2}{2}$ 44 $\frac{3}{2}$
LR-HD-A $\frac{3}{4}d$ $\frac{3}{4}d$ $\frac{d+\sigma}{2}$ $\frac{d+\sigma}{2}$ $\frac{2d}{d+\sigma}$
LR-HD-B $\frac{3}{4}d$ $\frac{9}{4}\sigma$ $\frac{d+\sigma}{2}$ $2\sigma$ $\frac{3}{2}$
Finally, we address several key boundaries and points. First, $d=4$ ($\sigma>2$) is the SR thermodynamic upper critical dimension, where logarithmic corrections to mean-field scaling arise in thermodynamic quantities. Similar logarithmic corrections are also manifested in geometric properties; for instance, $C_{1}\sim L^{3}(\ln L)^{1/4}$ and $C_{2}\sim L^{3}(\ln L)^{-1/4}$, where the latter exponent is conjectured from numerical evidence [fang_complete_2020]. Second, $d=6$ ($\sigma>2$) marks the onset of SR-HD-B scaling: FK clusters become percolation-like, and the number of spanning clusters diverges logarithmically [fang_geometric_2022]. Third, in the LR sector, the line $d=2\sigma$ plays the role of the LR thermodynamic upper critical dimension, while $d=3\sigma$ is its geometric analogue, separating LR-HD-A from LR-HD-B. Indeed, logarithmic corrections occur along the $d=2\sigma$ line, but whether they appear along the $d=\min(3\sigma,6)$ remains an open question.

### v2

@@ V Long-range O($n$) spin model
Figure 2: Universality diagram for the O($n$) models in the $(d,\sigma)$ plane. Seven regimes are identified: the no finite-temperature phase transition region (I), the short-range universality regimes (II and V), the long-range regimes (III–VI), and the complete-graph-like region (CG, VII) with distinct scaling behaviors. The boundary at $\sigma=2$ separates the SR and LR universality classes, while $\sigma=1$ marks a crossover between two nonclassical LR regimes.
@@
where $\mathbf{S}_{i}$ is an $n$-component unit vector at site $i$, $N=L^{d}$ is the total number of sites, and $r_{ij}$ denotes the minimal-image distance on the torus [20, 9, 1]. Here, $n=1,2,3$ corresponds to the Ising, XY, and Heisenberg models, respectively. In contrast to percolation, where the order parameter is purely geometric, O($n$) models possess internal spin degrees of freedom. The competition between spatial dimension, interaction range, and internal symmetry leads to rich critical behavior.
@@
In the SR limit $\sigma\to\infty$, the O($n$) models are governed by a $\phi^{4}$ field theory with upper critical dimension $d_{c}=4$. For $d=1$, all O($n$) models have $T_{c}=0$. In $d=2$, the Mermin–Wagner theorem forbids spontaneous breaking of continuous symmetry at any finite temperature, so the nature of the transition depends on $n$: the Ising model undergoes a continuous transition into a long-range ordered phase, the XY model exhibits a BKT transition into a quasi-long-range ordered phase, and the Heisenberg model has no finite-temperature transition but shows asymptotically free behavior at low temperatures [43, 34, 59]. For $d>d_{c}$, the SR Gaussian fixed point determines the critical scaling with exponents $y_{t}^{\rm GF}=2$ and $y_{h}^{\rm GF}=\frac{2+d}{2}$. In finite systems with PBCs, complete-graph asymptotics with $(y_{t}^{\rm CG},y_{h}^{\rm CG})=(\frac{d}{2},\frac{3d}{4})$ are required to describe the finite-size scaling fully, as discussed in Sec. II.
@@
When long-range interactions are introduced, the nature of the transition can change substantially, in ways that are parallel to LR-percolation but modified by the internal symmetry. In $d=1$, where SR Ising, XY, and Heisenberg models have no finite-temperature transition, sufficiently slowly decaying interactions ($\sigma\leq 1$) induce nontrivial criticality, and a BKT-like transition has been reported at $(d,\sigma)=(1,1)$ in the LR-Ising and LR-XY models [11, 49]. For $2\leq d<4$, numerical and field-theoretical analyses indicate three regimes as $\sigma$ is varied. For $\sigma>2$, the LR tail is irrelevant, and SR universality is retained. In a nonclassical regime $d/2<\sigma\leq 2$, the critical exponents become nontrivial functions of $\sigma$. Finally, for $0<\sigma\leq d/2$, the LR-GFP controls the transition, with exponents given by dimensional analysis as $(y_{t}^{\rm GF},y_{h}^{\rm GF})=(\sigma,(d+\sigma)/2)$. Extensive numerical evidence in $d=2$ for LR-Ising, XY, and Heisenberg systems supports a sharp SR–LR crossover at $\sigma_{*}=2$ [56, 58]. In particular, the point $(d,\sigma)=(2,2)$ already differs from the SR case: 2D LR-XY and LR-Heisenberg models exhibit second-order transitions, and the 2D LR-Ising exponents show detectable deviations from their SR values [56, 58, 57].
@@
Unlike the LR-percolation model, rigorous results for LR-O($n$) models in the low-dimensional nonclassical regime are lacking, but heuristic guidance can be obtained from LR-SRW. The LR-SRW effectively describes the Goldstone modes in the ordered phase of LR spin systems [57]. Specifically, when incorporating a fixed number of steps of $O(L^{d})$ to account for the extensivity of a finite system, the LR-SRW correctly reproduces the finite-size scaling of observables in the LRO phase, such as magnetic susceptibility. Notably, this framework succeeds even at the marginal case $\sigma=2$, where it correctly captures the existence of LRO and the logarithmic scaling of correlations that the standard Gaussian Free Field description fails to predict [56, 58, 57]. The LR-SRW also provides a simple geometric explanation for the no-transition regime, in analogy to the argument in the LR-percolation case. Given that the LR-SRW captures these essential physical features, the transition of the random walk from superdiffusive to hyperballistic at $\sigma=1$ leads us to conjecture a similar change in the critical behavior of O($n$) models at this boundary.
@@
Combining SR field theory, LR-GFP analysis, numerical results, and Lévy-flight arguments yields the universality diagram shown in Fig. 2. The overall topology in the $(d,\sigma)$ plane mirrors that of LR-percolation: we again distinguish seven regimes, labeled $T_{c}=0$, SR-LD, LR-LD-A, LR-LD-B, SR-HD, LR-HD, and CG. The main structural differences are that the SR upper critical dimension is now $d_{c}=4$ (instead of $6$ for percolation) and that the boundary between LR low-dimensional and LR high-dimensional regimes is given by $\sigma=d/2$. Since several sectors closely parallel the percolation case discussed in Sec. IV, here we only summarize the main regimes and emphasize features specific to O($n$) models.
@@
  • Regime VI (LR-HD): $d>2\sigma$ and $0<\sigma<2$. The LR Gaussian fixed point with $(y^{\rm GF}_{t},y^{\rm GF}_{h})=\left(\sigma,\frac{d+\sigma}{2}\right)$ controls the critical scaling in the thermodynamic limit, while the CG exponents $(y_{t}^{\rm CG},y_{h}^{\rm CG})=(\frac{d}{2},\frac{3d}{4})$ govern the finite-size scaling of global observables, as in the LR-HD regime of LR-percolation.
  • Regime VII (CG): $d\leq\sigma\leq 0$. The system exhibits the same critical scaling behaviors as the Ising model on the complete graph, governed by the exponents $(y_{t}^{\rm CG},y_{h}^{\rm CG})=(d/2,3d/4)$.
@@
  • $\sigma=d$ for $d<2$: This line marks the boundary of the $T_{c}=0$ regime, based on LR SRW arguments. As in LR-percolation, the nature of phase transitions along this line remains unknown, in general. The point $(d,\sigma)=(1,1)$ is a notable example where the LR-Ising model exhibits BKT-like transition.
@@
  • $\sigma=2$ for $2<d<4$: The line $\sigma=2$ separates SR and LR universality sectors but remains LR-dominated rather than SR-like. Numerical studies at $d=2$ show that at $(d,\sigma)=(2,2)$ the XY and Heisenberg models exhibit second-order transitions and the Ising exponents deviate from their SR values [56, 58, 57]. More generally, discontinuities in exponents and universal ratios as functions of $\sigma$ are expected when crossing this line within the LR-LD-A sector.
@@
  • $d=2\sigma$, $0<\sigma<2$: The LR analogue of the SR upper critical dimension. Here, the LR-GFP becomes marginal, and logarithmic corrections are expected, similarly to the role played by $d=4$ in the SR case.
@@ VI Long-range FK-Ising model
In addition to the standard spin representation, the Ising model admits a geometric formulation via the FK random-cluster representation [23], with the partition function defined in Eq. (1). In this framework, percolation and the Ising model share a common geometric language expressed through FK clusters.
Recent studies on the FK representation of the Ising model indicate the existence of two distinct upper critical dimensions [16, 17, 50, 12]. The standard thermodynamic upper critical dimension, $d_{c}=4$, marks the onset of mean-field critical behaviors. On the other hand, the cluster geometry exhibits a second, geometric upper critical dimension at $d_{p}=6$, above which the critical clusters, except for the largest one, become percolation-like. Following Ref. [16], it is convenient to parameterize the scaling of the largest and second-largest FK clusters by four fractal dimensions $(d_{\mathrm{L}1},d_{\mathrm{F}1},d_{\mathrm{L}2},d_{\mathrm{F}2})$ defined through
@@
where $C_{1}$ and $C_{2}$ are the sizes of the largest and second-largest FK clusters, $R_{1}$ and $R_{2}$ are their unwrapped gyration radii, and $L$ is the linear system size. Here, $d_{\mathrm{L}i}$ denotes the finite-size fractal dimension, characterizing the scaling with $L$, while $d_{\mathrm{F}i}$ represents the thermodynamic fractal dimension, relating the cluster size to its unwrapped gyration radius. For $4<d<6$, the FK clusters are geometrically Ising-like, described by $d_{\mathrm{F}1}=d_{\mathrm{L}1}=\frac{3d}{4}$ and $d_{\mathrm{F}2}=d_{\mathrm{L}2}=\frac{2+d}{2}$. The scaling of the largest cluster size is consistent with asymptotics derived rigorously on the CG Ising model [40, 6], while the scaling of the second-largest clusters is inferred from numerical results and found to be consistent with the GFP asymptotics [16, 17]. The winding number of the clusters and the number of spanning clusters are both of order $O(1)$. For $d\geq 6$, the finite-size fractal dimensions $d_{\mathrm{L}1}$ and $d_{\mathrm{L}2}$ remain the same as in the lower-dimensional case, but the thermodynamic fractal dimensions of clusters become $d_{\mathrm{F}1}=9/2$ and $d_{\mathrm{F}2}=4$, independent of dimensionality. Notably, $d_{\mathrm{F}2}=4$ corresponds to high-dimensional percolation; thus, the critical clusters, except for the largest one, show percolation-like behavior.
Moreover, the cluster number density $n(s,L)$ differs in the two regimes. Generally, one expects $n(s,L)\sim s^{-\tau}\tilde{n}(s/s_{c})$, where $s_{c}$ is a cutoff size close to the size of the largest cluster, scaling as $s_{c}\sim L^{d_{\mathrm{L}1}}$. The Fisher exponent $\tau$ is believed to satisfy the hyperscaling relation $\tau=1+d/d_{\mathrm{L}1}$, which has been confirmed for percolation in various dimensions and for the random-cluster model for $d<4$ [25]. For $4<d<6$, the Fisher exponent follows the GFP prediction as $\tau=1+d/y_{h}^{\rm GF}$ with $y_{h}^{\rm GF}=\frac{2+d}{2}$. However, for $d\geq 6$, the hyperscaling breaks down and one finds $\tau=5/2$, consistent with high-$d$ percolation. These differences in geometric scaling relations further lead to distinctive topological properties in the two regimes [16, 17].
@@
The same mechanism extends naturally to the LR-HD sector for $0<\sigma<\min(2,d/2)$, where the FSS of thermodynamic properties is governed by the LR Gaussian fixed point with $(y_{t}^{\rm GF},y_{h}^{\rm GF})=(\sigma,(d+\sigma)/2)$ and CG-Ising asymptotics. We conjecture that there exists a $\sigma$-dependent geometric upper critical dimension at $d_{p}=3\sigma$ within the LR mean-field regime, which divides it into two new regimes, LR-HD-A and LR-HD-B, as shown in Fig. 3.
  • Regime VI (LR-HD-A): $2\sigma<d<3\sigma$ with $0<\sigma<2$. Analogous to the SR-HD-A regime, the largest cluster follows the CG-Ising asymptotics, and the second largest clusters follow the LR-GFP scaling, i.e., $d_{\mathrm{F}1}=d_{\mathrm{L}1}=\frac{3d}{4}$ and $d_{\mathrm{F}2}=d_{\mathrm{L}2}=\frac{\sigma+d}{2}$. Both clusters scale linearly with the system size, as $R_{1}\sim R_{2}\sim L$. By analogy, we expect small clusters to have fractal dimension $d_{\mathrm{L}2}=(d+\sigma)/2$, and the Fisher exponents are $\tau=1+2d/(d+\sigma)$. Therefore, the number of spanning clusters remains finite in the thermodynamic limit.
  • Regime VIII (LR-HD-B): $d\geq 3\sigma$ and $0<\sigma<2$. In this regime, the critical clusters, except for the largest one, are expected to exhibit LR-percolation-like behavior, i.e., $d_{\mathrm{F}2}=2\sigma$, while the largest cluster is conjectured to have $d_{\mathrm{F}1}=\frac{9}{4}\sigma$. A notable difference from the SR-HD-B regime is that both $d_{\mathrm{F}1}$ and $d_{\mathrm{F}2}$ are $\sigma$-dependent, and the latter is consistent with those of the LR-percolation as listed in Table 2. The finite-size fractal dimensions $d_{\mathrm{L}1},d_{\mathrm{L}2}$ remain the same as in LR-HD-A. The unwrapped gyration radii of clusters are then given by,
@@
    Similar to the SR-HD-B regime, these clusters wind extensively across the periodic boundaries. Additionally, based on the insight of the CG percolation model [26], the Fisher exponent is also expected to be the same as the high-$d$ percolation case, with $\tau=5/2$. The number of spanning clusters then grows as $N_{s}\sim L^{d-3\sigma}$.
@@
This geometric upper critical dimension can be further corroborated by the structure of scaling windows near the critical point. As detailed in Ref. [16, 17] for the SR case, for $d>4$, the FK-Ising model exhibits two critical scaling windows. The first one is of size $O(L^{-d/2})$, consistent with the CG-Ising exponent $y_{t}^{\rm CG}=\frac{d}{2}$. In this interval, the FSS behaviors are identical to those at the critical point. However, a second scaling window exists in the high-T regime, where the FSS differs. For $4<d<6$, the size of the second window is controlled by the GFP exponent $y_{t}=2$. In this scaling window, the various observables follow the GFP scaling. For $d>6$, a percolation scaling window of size $O(L^{-d/3})$ emerges, where the scaling of all clusters is percolation-like. Similar scaling behaviors are expected in the LR-FK-Ising model and are summarized below.
@@
SR-HD-A  (V) $\frac{3}{4}d$ $\frac{3}{4}d$ $\frac{d+2}{2}$ $\frac{d+2}{2}$ $\frac{2d}{d+2}$
SR-HD-B  (VII) $\frac{3}{4}d$ $\frac{9}{2}$ $\frac{d+2}{2}$ 44 $\frac{3}{2}$
LR-HD-A  (VI) $\frac{3}{4}d$ $\frac{3}{4}d$ $\frac{d+\sigma}{2}$ $\frac{d+\sigma}{2}$ $\frac{2d}{d+\sigma}$
LR-HD-B  (VIII) $\frac{3}{4}d$ $\frac{9}{4}\sigma$ $\frac{d+\sigma}{2}$ $2\sigma$ $\frac{3}{2}$
Finally, we address several key boundaries and points. First, $d=4$ ($\sigma>2$) is the SR thermodynamic upper critical dimension, where logarithmic corrections to mean-field scaling arise in thermodynamic quantities. Similar logarithmic corrections are also manifested in geometric properties; for instance, $C_{1}\sim L^{3}(\ln L)^{1/4}$ and $C_{2}\sim L^{3}(\ln L)^{-1/4}$, where the latter exponent is conjectured from numerical evidence [15]. Second, $d=6$ ($\sigma>2$) marks the onset of SR-HD-B scaling: FK clusters become percolation-like, and the number of spanning clusters diverges logarithmically [16]. Third, in the LR sector, the line $d=2\sigma$ plays the role of the LR thermodynamic upper critical dimension, while $d=3\sigma$ is its geometric analogue, separating LR-HD-A from LR-HD-B. Indeed, logarithmic corrections occur along the $d=2\sigma$ line, but whether they appear along the $d=\min(3\sigma,6)$ remains an open question.